A company issues 9%, 7-year bonds with a par value of $260,000 on January 1 at a price of $273,732, when the market rate of interest was 8%. The bonds pay interest semiannually. The amount of each semiannual interest payment is:

A) $23,400. B) $11,700. C) $0. D) $20,800. E)$10,400.

A company must repay the bank a single payment of $26,000 cash in 6 years for a loan it entered into. The loan is at 7% interest compounded annually. The present value factor for 6 years at 7% is .6663. The present value of an annuity factor for 6 years at 7% is 4.7665.The present value of the loan (rounded) is:

A) $26,000. B) $123,929. C) $17,324. D) $5,455. E) $23,086.

On July 1, Shady Creek Resort borrowed $400,000 cash by signing a 10-year, 9% installment note requiring equal payments each June 30 of $62,328. What is the journal entry to record the first annual payment?

A) Debit Cash $400,000; debit Interest Expense $62,328; credit Notes Payable $462,328.

B) Debit Interest Expense $62,328; credit Cash $62,328.

C) Debit Interest Expense $36,000; credit Cash $36,000.

D)Debit Interest Expense $36,000; debit Notes Payable $26,328; credit Cash $62,328.

E) Debit Interest Expense $36,000; debit Interest Payable $26,328; credit Cash $62,328

A company purchased equipment and signed a 5-year installment loan at 8% annual interest. The annual payments equal $10,400. The present value of an annuity factor for 5 years at 8% is 3.9927. The present value of a single sum factor for 5 years at 8% is .6806. The present value of the loan is:

A) $7,078. B) $52,000. C) $41,524. D) $10,400. E) $15,281.

On January 1, Parson Freight Company issues 7.5%, 10-year bonds with a par value of $2,600,000. The bonds pay interest semiannually. The market rate of interest is 8.5% and the bond selling price was $2,423,327. The bond issuance should be recorded as:

A) Debit Cash $2,423,327; credit Bonds Payable $2,423,327.

B) Debit Cash $2,600,000; credit Bonds Payable $2,600,000.

C) Debit Cash $2,423,327; debit Interest Expense $176,673; credit Bonds Payable $2,600,000.

D) Debit Cash $2,600,000; credit Bonds Payable $2,423,327; credit Discount on Bonds Payable $176,673.

E) Debit Cash $2,423,327; debit Discount on Bonds Payable $176,673; credit Bonds Payable $2,600,000.
Answer #1

1) A company issues 9%, 7-year bonds with a par value of $260,000 on January 1 at a price of $273,732, when the market rate of interest was 8%. The bonds pay interest semiannually. The amount of each semiannual interest payment is:

Interest payment = 260000*9%*6/12 = $11700

So answer is b) $11700

2) A company must repay the bank a single payment of $26,000 cash in 6 years for a loan it entered into. The loan is at 7% interest compounded annually. The present value factor for 6 years at 7% is .6663. The present value of an annuity factor for 6 years at 7% is 4.7665.The present value of the loan (rounded) is:

Present value = (26000*.6663) = $17323.80 or $17324

So answer is c) $17324

3) On July 1, Shady Creek Resort borrowed $400,000 cash by signing a 10-year, 9% installment note requiring equal payments each June 30 of $62,328. What is the journal entry to record the first annual payment?

Interest expense = 400000*9%* = 36000

So answer is D)Debit Interest Expense $36,000; debit Notes Payable $26,328; credit Cash $62,328

4) A company purchased equipment and signed a 5-year installment loan at 8% annual interest. The annual payments equal $10,400. The present value of an annuity factor for 5 years at 8% is 3.9927. The present value of a single sum factor for 5 years at 8% is .6806. The present value of the loan is:

Present value of loan = 10400*3.9927 = 41524

So answer is c) $41524

5) On January 1, Parson Freight Company issues 7.5%, 10-year bonds with a par value of $2,600,000. The bonds pay interest semiannually. The market rate of interest is 8.5% and the bond selling price was $2,423,327. The bond issuance should be recorded as:

So answer is E) Debit Cash $2,423,327; debit Discount on Bonds Payable $176,673; credit Bonds Payable $2,600,000.